OpenStudy (anonymous):

What is the equation of the line that passes through the point (2, 3) and is perpendicular to the line y = –x – 3?

OpenStudy (anonymous):

Equation of the line is y=-x-3 Its slope m=-1 Slope of line perpendicular to above line=-1/m=-1/-1=1 Equation of line with slope 1 and passing through (2,3) is y-3=1(x-2) y=x-2+3 Required eq. is y=x+1

OpenStudy (anonymous):

thanks so much